A 45ft wire is split into 3 pieces. One is 15 ft,another is to be 2 ft less than 3 times the length of the last? Find the length of all the pieces?
Answer by mananth(16946)

one is 15 ft.
let the third be x feet
the second is 3x-2.
...
15+x+3x-2 = 45
13+4x=45
add -13 to both sides
13-13+4x=45-13
4x=32
divide by 4
x= 32/4
x= 8 one piece.
second piece = 3x-2 = 3*8-2
= 22
..
15,8,22 . add them up you will get 45